KOFI electronics bought a shipment of TVs at a net price of $477.36 each, after discounts of 15%, 10%, and 4%. What is the list price?


1
Expert's answer
2021-03-09T16:45:12-0500

let "x" represent the price of each items and A, B, C represent each item

.: "price -discount =netprice"

i) item A:

"x-15x"% = 477.36

"x-\frac{15x}{100}=477.36"

"\frac{85x}{100}=477.36"

"85x" = 100 × 477.36

"x" = "\frac{100 × 477.36}{85}"

"x=561.6"

ii) item B:

"x-10x"% = 477.36

"x-\frac{10x}{100}=477.36"

"90x=" 100 × 477.36

"x=\frac{100 × 477.36}{90}"

"x=530.4"

iii) item C:

"x-4x"% = 477.36

"x-\frac{4x}{100}=477.36"

"96x=100\u00d7477.36"

"x=\frac{100 × 477.36}{96}"

"x=497.25"


Therefore, the list price of each of the item is $561.6, $530.4 and $497.25